### v1

# Optimizing Visible Light Communication Efficiency Through Reinforcement Learning-Based NOMA-CSK Integration

Abstract
In this paper, we explore the use of Non-Orthogonal Multiple Access (NOMA) and Color Shift Keying (CSK) for Visible Light Communication (VLC) systems. VLC is a wireless communication technology that uses visible light as the carrier signal to transmit information. It has several advantages over traditional radio frequency communication, including higher bandwidth, lower interference, and greater security. We first provide an introduction to NOMA and CSK and explain how they can be applied to VLC systems. NOMA is a technique that allows multiple users to share the same frequency channel by allocating different power levels to each user. This enables more users to connect to a single VLC transmitter simultaneously, thereby improving system capacity and spectral efficiency. CSK, on the other hand, is a modulation technique that uses different colors of light to represent digital information. By changing the color of the transmitted signal, information can be encoded and decoded at the receiver. Next, we discuss how NOMA and CSK can be combined in VLC systems by using different power levels to represent different users. This allows for more efficient use of the frequency spectrum, as multiple users can share the same channel at the same time. Additionally, we examine the potential benefits of using NOMA and CSK together in VLC systems to increase data rate. Finally, we discuss how reinforcement learning, a machine learning technique used to train agents to make decisions based on environmental feedback, can be used to optimize NOMA-CSK-VLC networks by allowing agents to learn and adapt to changing network conditions. Overall, our paper provides insights into the benefits of combining NOMA and CSK for VLC systems, highlighting the potential for improving communication efficiency and performance.
keywords:
NOMA, CSK, power allocation, RL, SACAgent
1 Introduction
With the development and widespread use of wireless communication systems, the number of wireless devices and data transfer requirements are increasing. However, the current frequency band of 6 GHz and below has become too crowded to provide efficient communication [1]. Therefore, it is suggested to use wide unlicensed spectrum of the visible light in wireless communication systems [2]. Visible light communication (VLC) systems offer a number of benefits, including energy savings due to the ability to carry out lighting and communication simultaneously, security, privacy, and resistance to electromagnetic radiation [3, 4].
VLC systems usually utilize light emitting diodes (LEDs) as transmitters, while photodetectors (PDs) are favored as receivers due to their quick response times. The modulation speeds of LEDs are a crucial element in communication performance. While phosphor-coated white LEDs have modulation speeds of approximately 10 MHz, multicolored LEDs, such as Red, Green Blue, (RGB) and Quadrature LED (QLED), can be modulated at speeds of almost 100 MHz. Therefore, compared to white LEDs with phosphor coating, communication via multicolored LEDs may theoretically be possible at higher speeds.
In wireless communication systems, modulation type, which is one of the factors affecting communication performance, is as important as date rates. Various modulation methods such as On-Off Keying (OOK) [5], Pulse Width Modulation (PWM) [6], Color Shift Keying (CSK) [2] can be used as modulation type in VLC systems. The newest of these, CSK, is a modulation method published in the IEEE 802.15.7 standard and used in VLC systems with multi-colored LED transmitters [2]. The CSK modulation method is a modulation method in which the color of the light emitted from the transmitter is changed according to the information signal in multicolor LED VLC systems. On the receiver side, there are color filters. Signals obtained from color filters are used in demodulation. For RGB LED emitter systems, the IEEE 802.15.7 standard defines 4-CSK, 8-CSK, and 16-CSK.
CSK has several advantages in VLC systems. Firstly, it maintains a constant power envelope of the transmitted signal, which reduces the possibility of health issues caused by fluctuations in light intensity. Secondly, it provides color lights with a stable temperature. Thirdly, with rapid switching, the flicker effect is minimized. Fourthly, even at slow switching speeds, high-level modulation is possible, which increases the capacity [2]. Additionally, CSK does not require additional DC power for data transmission, giving it an edge over other modulation schemes utilized in VLC. This allows for an increase in the signal-to-noise ratio, and efficient use of the entire transmission spectrum is possible.
CSK modulation involves instantaneously changing the color of light to transmit data. As a result, lighting characteristics such as Correlated Color Temperature (CCT) [7], Color Rendering Index (CRI) [8], and illuminance [8] can also change during communication. CCT refers to the temperature of the light in Kelvin in the given environment, while CRI is a measure of how well the light source can accurately reflect the colors of objects in the environment. CRI is typically represented as a value between 0 and 100, with higher values indicating more accurate color representation.
Besides CSK modulation, multiple access methods in VLC have also been extensively studied in the literature. Both orthogonal and non-orthogonal multiple access methods (OMA and NOMA) are available in VLC. Power domain NOMA (PD-NOMA), which is a non-orthogonal multiple access method, has been proposed as a promising and bandwidth-efficient solution for multiple access in indoor VLC networks [9, 10]. NOMA works by superimposing user signals in the power domain, where a separate power level is assigned to each signal based on the user’s channel conditions. In NOMA, all users have continuous access to full bandwidth. The fundamental tenet of NOMA is to give users with weak channels more power than those with strong channels. To achieve this, the user with the highest power can detect the information signal despite the interference caused by other users, while other users perform successive interference cancellation (SIC) to decode their own signals [11]. It has been demonstrated in [12] that NOMA can achieve higher data rates than Orthogonal Frequency Division Multiple Access (OFDMA) in VLC systems .Similarly, in [13] researchers examined the performance of indoor VLC systems using NOMA and discovered a substantial enhancement in system capacity due to the implementation of NOMA. In [14], NOMA techniques were compared with classical OFDMA techniques, and it was found that some of the advantages of NOMA include better quality of service, less interferences, and the ability to provide multiple access to a larger number of users.
Power allocation is a crucial aspect of NOMA VLC systems in determining the achievable capacity [15]. The gain ratio power allocation (GRPA) method was proposed in [9] for a scenario with two transmitters and three receivers, and it showed better sum-rate and Bit Error Rate (BER) than static power allocation in the best-case scenario. In [16], the normalized gain difference power allocation (NGDPA) method was introduced and compared to GRPA, resulting in 29.1% higher total capacity in a 3-user 2x2 MIMO VLC system. An optimization algorithm was proposed in [17] to maximize overall data rate and compared to GRPA. However, fairness and Quality of Service (QoS) were not considered. In [18], a convex decoder was used to optimize the total data rate and BER under a certain light intensity constraint, but users’ data rate and fairness were not considered. The VOOK method with GRPA was used in [5] for power allocation in NOMA-VLC systems with variable illumination levels. However, VOOK is inefficient as the total capacity decreases in proportion to brightness. In [19], a particle swarm optimization (PSO)-based power allocation optimization algorithm for NOMA VLC multicellular networks was proposed, but the level of fairness achieved was less than 0.6, which is suboptimal for QoS. When distributing LED transmitter power to users in VLC systems, factors such as fairness, illumination level, and user data rate should be taken into account to provide eye comfort and better QoS.
Recent researches have explored the use of reinforcement learning (RL) methods to optimize communication systems. Using deep reinforcement learning (DRL) as the method, Xiao et al. [20] investigated power allocation for non-orthogonal multiple access (NOMA) against smart jammers in radio frequency (RF) systems, resulting in an increased sum-rate. He et al. [21], on the other hand, studied power distribution in multi-carrier RF systems with NOMA and observed an overall system performance improvement. Zhang et al. [22] employed DRL to investigate energy-efficient resource allocation, while Zhang et al. [23] demonstrated a 32.9% higher efficiency compared to the ALOHA method. Wang et al. [24] proposed the DRL-JRM (Deep Reinforcement Learning-Joint Resource Management) scheme for RF multi-carrier systems in NOMA, which showed superior system efficiency and resistance to interference, especially in the presence of large numbers of users and strong intercellular interference. Bariah et al. [25] investigated Deep Q-Learning-Based resource allocation to increase the energy efficiency and sum-rate for NOMA VLC systems, showing its superiority over the results obtained by genetic algorithm and differential evolution. Finally, Al et al. [26] investigated Q-Learning-Based resource allocation and LED emitter angle optimization to increase sum-rate for NOMA VLC systems.
In this study, a PD-NOMA scheme was implemented in a VLC system using CSK modulation. Reinforcement learning (RL) with a Soft Actor-Critic (SAC) agent [27] was trained to maximize fairness in power allocation. The agent was trained to be applicable not only to a specific indoor environment, but to other indoor environments as well. The results showed that the proposed RL-based PD-NOMA CSK VLC system outperformed other methods in terms of fairness, sum-rate and bit error rate. Moreover, the trained agent was able to achieve the same performance in different indoor environments for the CSK VLC system without the need for retraining.
2 VLC Channel
The VLC channel can be modeled in two ways: line of sight (LoS) and non-line of sight (NLoS). According to [28], the strongest NLoS channel of the VLC channel is 7dB weaker than the weakest LoS channel. Therefore, this study is based on the LoS model. According to the model in [29], the attenuation of light is established for the case of direct vision between the transmitter and receiver and its mathematical expression is given in Eq.1.
$H_{LoS}=\frac{A_{r}R\left(\phi_{T_{x}}\right)}{d^{2}}T_{S}(\psi)g(\psi)\cos(\psi)$ (1)
where $R(\phi_{Tx})=\frac{(k_{l}+1)}{2\pi}\cos^{k_{l}}(\phi_{Tx})$ is the LED’s radiant intensity, $k_{l}=(-\ln 2)/\ln(\cos(\Omega_{1/2}))$ is the Lambertian emission mode number, $d$ is the distance between transmitter and receiver, $\phi_{Tx}$ is the radiation angle of the LED, $\Omega_{1/2}$ is the LED half angle, $A_{r}$ is the surface area of the photodiode (PD), $\psi$ is the incidence angle of the PD, $T_{s}(\psi)$ is the optical filter gain, $g(\psi)=n^{2}/\sin^{2}\phi$ is the optical lens gain, $n$ is the refractive index of the PD, and $\phi$ is the view angle of the PD.
3 System Model
For the simulation purposes, an office space measuring 5 meters in length, 5 meters in width, and 3 meters in height will be utilized. The OSRAM LZ4-00MA00 multicolor LED, which is commercially available, has been selected for this particular application. The assumption made is that the emitted lighting powers for the red (R), green (G), and blue (B) colors are equivalent.
LEDs can be represented in various ways for simulation purposes, and the H model proposed in [30] closely approximates the results obtained from experimental studies. The H model parameters for a specific LED can be found in Table 1 [31], and these parameters are utilized to compute the values for Eq.2, Eq.3 and Eq.4.
$g\left(\lambda,\lambda_{p},\Delta\lambda\right)=\exp{\left[-\frac{\left(\lambda-\lambda_{p}\right)^{2}}{\mathrm{\Delta}\lambda^{2}}\right]}$ (2)
$\Delta\lambda=\left\{\begin{matrix}\Delta\lambda_{1},&\lambda<\lambda_{p}\\ \Delta\lambda_{2},&\lambda\geq\lambda_{p}\\ \end{matrix}\right.$ (3)
$PSD_{i}\left(\lambda\right)=\frac{g\left(\lambda,\lambda_{p},\Delta\lambda\right)+k_{1}g^{k_{2}}\left(\lambda,\lambda_{p},\Delta\lambda\right)}{1+k_{1}}$ (4)
Here, $\lambda_{p}$ is the peak wavelength in the $PSD_{i}$ while $i$ indicating the color, $\Delta\lambda_{1}$ and $\Delta\lambda_{2}$ are the left and right half spectral widths, respectively, and $k_{1}$ and $k_{2}$ are characteristic parameters of the PSD shape. The $PSD\left(\lambda\right)$ for each color obtained from this H model is shown in Fig.0(a).Thirty-six of these LEDs were placed on the ceiling of the office, at a distance of 95 cm from the side walls in 4 square luminaires with a distance of 3 cm between each LED, as in Fig.0(b). X and Y denote the coordinates of the bottom-left position referenced room.
Table 1: Parameter values for different color components [31]
Parameter Red Green Blue
$\lambda_{p}$ 632.5 517.7 453
$\Delta\lambda_{1}$ 23.84 29.38 18.99
$\Delta\lambda_{2}$ 14.74 45.21 25.5
$k_{1}$ 2 2 2
$k_{2}$ 6 3 5
(a) Power spectra of the LZ4-00MA00 without amber color
(b) Locations of LED transmitters
Figure 1: Power spectra and LED transmitter locations
The R, G, and B spectra’s intensity changes when the data to be transmitted is modulated using CSK modulation. Color filters can be used at the receiver to identify changes in these intensities. In this instance, Eq.5 describes the power of the signals that were received from the $j$-th color filter.
[…]
In the context provided, $i$ and $j$ correspond to the colors of the LED at the transmitter and the photodetector at the receiver, respectively. The values of $i$ and $j$ range from $1$ to $3$, representing the colors red, green, and blue. The variables $\lambda_{l}^{i}$ and $\lambda_{h}^{i}$ denote the lower and upper wavelengths of the spectrum for the LED color $i$, respectively. The variable $F_{j}$ represents the gain associated with the $j$-colored optical filter. The variable $S_{i}$ denotes the CSK symbol value for color $i$. The matrix $\boldsymbol{p}$ is a $3\times 3$ channel matrix, where the $(i,j)$-th element is determined by Eq. 6.
$\boldsymbol{p}=\left[\begin{matrix}p_{1,1}&p_{2,1}&p_{3,1}\\ p_{1,2}&p_{2,2}&p_{3,2}\\ p_{1,3}&p_{2,3}&p_{3,3}\\ \end{matrix}\right]$ (6)
The non-ideal nature of both the optical filters and the LEDs results in the matrix $\boldsymbol{p}$ not being a diagonal matrix. The elements outside the diagonal reflect the power of interference. In this case, the SINR at the receiver is:
$\operatorname{SINR}=\frac{\sum\limits_{i}\sum\limits_{j}p_{i,j}\delta_{ij}}{\sum\boldsymbol{p}-\sum\limits_{i}\sum\limits_{j}p_{i,j}\delta_{ij}+N_{0}}$ (7)
Here, $\delta_{ij}$ is Kronecker delta function and $N_{0}$ is the power of AWGN noise respectively. Capacity calculation over SINR can be made done according to Eq.8. After the CSK demodulation is performed on the received signal, the throughput can be calculated using Eq.9.
$C=Blog_{2}\left({1+SINR}\right)$ (8)
$T=\ C\left(1-BER\right)$ (9)
The open-source program shared in [32] can be used to calculate the CRI, CCT, and illumination parameters mentioned in the introduction. The irradiance data corresponding to the LED’s wavelength is input into the program, which then uses it to calculate the illumination parameters. The CRI, CCT, and illumination values should be calculated over the long-term average of irradiance value because the brightness of the light emitted from the R, G, and B colors varies with communication.
4 Simulations
The variables associated with the VLC channel and their corresponding values, along with the simulation parameters, are provided in Table 2. We built a NOMA-CSK communication system with two receivers ($U_{1}$ and $U_{2}$). Both receivers’ PDs are positioned with their normal surfaces perpendicular to the ceiling.
Table 2: VLC channel and simulations parameters
Parameter Value
RGB LED viewing angle $60^{\circ}$
Power of an RGB LED 1 W
PD responsivity 0.54
PD surface area $1~{}\mathrm{cm}^{2}$
Optical ideal filter lower cutoff wavelengths [590, 485, 380] [R,G,B] nm
Optical ideal filter upper cutoff wavelengths [764, 590, 485] [R,G,B] nm
Optical filter gain 1
Optical lens gain 1
Speed of light 299,792,458 m/s
Transmitter-receiver bandwidth 30 MHz
SNR 30 dB
Mudolation Level 4
Both users have three PDs with suitable color filters in front of each PD for the VLC-NOMA-CSK simulation. The following Eq.10 describes how NOMA is applied to the CSK VLC system:
[…]
Here, $S_{i}^{(1)}$ and $S_{i}^{(2)}$ are the value of the CSK symbol in color $i$ for the $U_{1}$ and $U_{2}$, respectively. $\rho^{(1)}$ is the NOMA power allocation coefficient for $U_{1}$, who has a bigger channel gain. ${{h^{(k)}}}$ is the total channel between tranmister and $k-th$ user. In this case, the SINR at the receivers when demodulation is performed with the help of SIC is:
$\operatorname{SINR}^{(1)}=\frac{\sum\limits_{i}\sum\limits_{j}p_{i,j}^{(1)}\delta_{ij}}{\sum\limits_{i}\sum\limits_{j}{({p_{i,j}}-p_{i,j}^{(1)}\delta_{ij})}+N_{0}^{(1)}}$ (11)
$\operatorname{SINR}^{(2)}=\frac{\sum\limits_{i}\sum\limits_{j}p_{i,j}^{(2)}\delta_{ij}}{\sum\limits_{i}\sum\limits_{j}{({p_{i,j}}-p_{i,j}^{(2)}\delta_{ij})}+N_{0}^{(2)}+N^{(1)}}$ (12)
$N_{0}^{(k)}=\frac{\left(\sum\limits_{i}\sum\limits_{j}\left|p_{i,j}^{(k)}\delta_{ij}\right|\right)^{2}}{SNR}$ (13)
$N_{0}^{(1)}$ and $N_{0}^{(2)}$ represent the AWGN noise powers for $U_{1}$ and $U_{2}$ respectively. $N^{(1)}$ denotes the interference caused by the message signals of $U_{1}$ on $U_{2}$, and it can be expressed as follows:
$N^{(1)}=\sum_{i}\sum_{j}\left(\int_{\lambda_{l}^{i}}^{\lambda_{h}^{i}}PSD(\lambda)F_{j}S_{i}^{(1)^{2}}p^{(1)}\left(h^{(1)}\right)^{2}\mathrm{~{}d}\lambda\right)$ (14)
In a practical communication system, the positions of users are not fixed, and the channel conditions change as they move. These factors have been considered in the proposed reinforcement learning-based NOMA-CSK power allocation method. Our objective was to train the agent to efficiently allocate power even when confronted with channel conditions that differ from those encountered during training. Reinforcement learning is a method for training an agent in a simulation environment through trial and error, without relying on a pre-existing dataset. During these trials, the agent learns a policy that maximizes rewards by receiving a positive reward when it achieves a desired outcome, and a negative reward (penalty) otherwise. In this study, a model-free, online, off-policy, actor-critic reinforcement learning agent called soft actor critic (SACA) was utilized. Since both the action and observation values are continuous, an agent capable of working in this environment is selected as the SACA [27].
The environment should be designed before the learning process can be implemented. The actions that the agent can take and the quantities it can observe should be determined to design the environment. The action to be taken by the agent is defined as changing the power allocation $\rho$ of the user with the highest channel gain within the range of [0 0.5]. In this case, the remaining power will be used by other users. The value that the agent will observe is the ratio of the channel gains of $U_{1}$ and $U_{2}$ to each other is given in Eq.15, and it is randomly determined.
$\mathrm{r}=\frac{h^{(2)}}{h^{(1)}},\mathrm{r}\in\left[\begin{array}[]{ll}0&1\end{array}\right]$ (15)
The users’ channel gains are determined to fall within the range of $2.84\times 10^{-5}$ and $5.98\times 10^{-4}$. This range encompasses the maximum and minimum values among 2601 channels, obtained by varying the users’ coordinates as $4.50.5<X<$ 4.5, $4.50.5<Y<$ 4.5, and $2.50.5<Z<$ 2.5, with an increment of 0.25 for each step. The values of $h^{(1)}$ and $h^{(2)}$ are generated randomly according to Eq.16 and Eq.17, respectively.
$2.84\times{10}^{-5}<{{h^{(1)}}}<\frac{5.98\times{10}^{-4}}{r}$ (16)
${{h^{(2)}}}={{h^{(1)}}}\times r$ (17)
After creating the environment, the agent’s performance criteria and reward mechanism should be determined. The agent will be trained to maximize fairness as an important objective in multiple access systems. Fairness is a measure of how fairly a system resource is shared among users or applications used in communication systems and is given in Eq.18.
$\frac{\left(\sum\limits_{k=1}^{N}T_{k}\right)^{2}}{N\sum\limits_{k=1}^{N}T_{k}^{2}}$ (18)
Here, $N$ represents the total number of users in the system. $T_{k}$ represents the throughput of the $k-th$ user. If the fairness value that can be calculated using Eq.18 is above 0.99, the agent is rewarded with 1; otherwise, it is rewarded with -10. The SACAgent is trained for 1000 episodes by randomly generating user channel ratios and channels. Training is completed when the average reward value of the last 20 episodes is 1. The agent is tested with 100 randomly generated pairs of $({{h^{(1)}}},{{h^{(2)}}})$ where the value of $r$ is varied in increments of 0.01 within the range of 0 to 1.
The changes in sum-rate, throughput, and fairness values are depicted in Fig.1(a), Fig.1(b), and Fig.1(c) as the channels ratio $r$ varies. These values were obtained by applying proposed power allocation techniques to users in randomly generated $({{h^{(1)}}},{{h^{(2)}}})$ channels. The figures also include the performance results of TDMA-based multiple access, GRPA and NGDPA. Upon examination, it becomes clear that the trained agent demonstrates significantly superior performance in sum-rate, throughput, and fairness compared to other methods.
(a) Sum-rate performance
(b) Throughput performance
(c) Fairness performance
(d) BER performance
Figure 2: Performance comparison of the agent, grpa, ngdpa and tdma
For a 30 dB SNR, a bit error rate (BER) of 0 was achieved for all channel rates in TDMA and AGENT, while the BER graphs obtained for GRPA and NGDPA are shown in Fig.1(d).
The PSD of the LED’s light output is altered due to modulation. In order to determine CRI , CCT, and illumination level, the modified PSD is used for calculations. Eq. 19 represents the power of the transmitted symbols of $U_{1}$ and $U_{2}$ in color $i$. By this, altered PSD ($PSD_{a}(\lambda)$) can be calculated using Eq. 20
$TSP_{i}=\mathbb{E}\left(\left[S_{i}^{(1)}\sqrt{p^{(1)}}+S_{i}^{(2)}\sqrt{1-p^{(1)}}\right]^{2}\right)$ (19)
$PSD_{a}(\lambda)=\sum\limits_{i=1}^{3}PSD_{i}(\lambda)TSP_{i}$ (20)
In this case, after obtaining $PSD_{a}$, CRI, CCT and illumination parameters for the cases with and without communication are calculated and given in Table 3. The NOMA method did not change the CRI, CCT and illumination level. Only a decrease in illumination value was observed due to modulation in communication.
Table 3: Illuminations parameters comparision table
CRI CCT Illumination (lux)
W/O Comm. 32.3121 9445 863.64
W Comm. 32.3121 9445 287.59
The agent was trained for a specific channel range. To investigate its performance in an environment other than where it was trained, we tested it in a new environment with a lower bound value 5 times smaller and an upper bound value 5 times larger than that of the environment it was trained in. We compared the agent’s performance with other power allocation methods. As can be seen from Fig.2(a), Fig.2(b), and Fig.2(c), the agent performs better in terms of sum-rate, throughput, and fairness compared to other methods in an environment outside of where it was trained. The BER graph for NGDPA and GRPA is given in Fig.2(d) while 0 BER is achieved with AGENT and TDMA.
(a) Sum-rate performance
(b) Throughput performance
(c) Fairness performance
(d) BER performance
Figure 3: Performance comparison of the agent, grpa, ngdpa and tdma in a different environment
5 Conclusion
In this paper, we explored the combination of Non-Orthogonal Multiple Access (NOMA) and Color Shift Keying (CSK) for Visible Light Communication (VLC) systems. The integration of NOMA and CSK was discussed, showcasing how power levels can represent different users and facilitate more efficient channel sharing. This synergy holds the promise of increasing data rates and improving communication efficiency in VLC systems. Additionally, we explored the application of reinforcement learning, specifically Soft Actor-Critic (SAC), to optimize NOMA-CSK-VLC networks. The agent learned to adapt power allocation based on changing network conditions, achieving high fairness and system performance.
Simulations were conducted to evaluate the proposed approach’s effectiveness in terms of sum-rate, throughput, fairness, and Bit Error Rate (BER). The results demonstrated that the reinforcement learning-based NOMA-CSK approach outperformed traditional power allocation methods in various environments, showcasing its robustness and adaptability.
In conclusion, our study sheds light on the potential benefits of combining NOMA and CSK in VLC systems, offering insights into enhanced communication efficiency and performance. The reinforcement learning approach presented a promising solution for optimizing power allocation in dynamic network scenarios. As VLC continues to evolve, the integration of advanced techniques like NOMA and CSK, along with intelligent optimization strategies, could pave the way for more efficient and reliable wireless communication systems.
Data availability Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
Declarations
Conflict of interest
We declare that we do not have any commercial or associative interest that represents a conflict of interest in connection with the work submitted.